OpenStudy (anonymous):

solve 2x^2+9x=−5 using the quadratic formula, round to the nearest hundred if more then one answer separate with commas.

OpenStudy (anonymous):

@hero

OpenStudy (johnweldon1993):

Well first...you need to set this equal to 0...so add 5 to both sides of the equation

OpenStudy (anonymous):

2,9,-5?

OpenStudy (anonymous):

5+2X^2+9X=-5+5 Like that?

OpenStudy (johnweldon1993):

Yes, just remember to order it correctly afterwards, So you would have \[\large 2x^2 + 9x + 5 = 0\] NOW what are your 'a' 'b' and 'c' ?

OpenStudy (anonymous):

a=2 b=9 c=5

OpenStudy (johnweldon1993):

Perfect! *just remember the equation always need to be = 0" So we have a = 2 b= 9 c = 5 The quadratic formula says \[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] So when we plug in all the numbers we have?

OpenStudy (anonymous):

-9+sqrt(9^2-4(2)(5)/2(2)

OpenStudy (anonymous):

\[-9+\sqrt{9^2-4(2)(5)}/2(2)\]

OpenStudy (johnweldon1993):

Mmhmm \[\large \frac{-9 \pm \sqrt{9^2 - 4(2)(5)}}{2(2)}\] so we simplify a little to \[\large \frac{-9 \pm \sqrt{81 - 40}}{4}\] what does the square root simplify to?

OpenStudy (anonymous):

\[\sqrt{41}\]

OpenStudy (johnweldon1993):

Right, so now we break it into our 2 equations \(\large \frac{-9 + \sqrt{41}}{4}\) and \(\large \frac{-9 - \sqrt{41}}{4}\) and just solve those 2...those are your 2 values

OpenStudy (anonymous):

does it go into a decimal?

OpenStudy (johnweldon1993):

it will go into a decimal yes...unless you're supposed to keep it exact in which case this will be done

OpenStudy (anonymous):

i need to round it to the nearest hundred

OpenStudy (johnweldon1993):

Ahh, then there you go, so evaluate the decimals and round to the nearest hundredth...and then you're all set :)

OpenStudy (anonymous):

1.−0.64921894 2.−3.85078105 idk how to round to the nearest hundred :(

OpenStudy (johnweldon1993):

That's okay Just remember  that's the first one...so rounded to the nearest hundredth would be -0.65 right?